The sum of Danny's and Tracy's ages is 27. Danny's age five years from now will be twice what Tracy's age was two years ago. How old is each person now?
Answer by sachi(548)

let danny is x yrs , so tracy is 27-x yrs
as per question x+5=2(27-x-2)
or x+5=50-2x
or x+2x=3x=50-5=45
or x= 45/3=15
so danny is 15 yrs & tracy is 12 yrs
ans